OpenStudy (anonymous):

Partial derivatives and chain rule problem: Show that\[z _{x}x _{y}y _{z}=-1\] given F(x,y,z)=0 and z=f(x,y) y=g(x,z) x=h(y,z) and \[F _{x} \neq 0 F _{y} \neq 0 F _{z} \neq 0\]

OpenStudy (anonymous):

Show that\[z _{x}x _{y}y _{z}=-1\] given F(x,y,z)=0 and z=f(x,y) y=g(x,z) x=h(y,z) and \[F _{x} \neq 0 F _{y} \neq 0 F _{z} \neq 0\]

OpenStudy (anonymous):

So why aren't you using chain rule?

OpenStudy (anonymous):

I believe I'm supposed to expand on the fact that dy/dx=-Fx/Fy Id like to simply say that Zx=-Fx/Fz (and so on) but I don't think that follows, and I'm not sure how to show that it doesn't

OpenStudy (anonymous):

\[ \frac{d}{dx}F(x,y,z) = F_x+F_yy_x+F_zz_x\\ \frac{d}{dy}F(x,y,z) = F_yx_y+F_y+F_zz_y\\ \frac{d}{dz}F(x,y,z) = F_xx_z+F_yy_z+F_z \]

OpenStudy (anonymous):

All of these equations should equal each \(0\) as well.

OpenStudy (anonymous):

So you'd say \[ z_x=-\frac{F_x+F_yy_x}{F_z} \]

OpenStudy (anonymous):

I got those equations for Xy Yz and Zx and multiplied them out to see if I could get -1 didn't

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

FyYzXy+1+FyYx+FxXz+ZyYzXz+FzZy+FzYxZy+FxXzZy

OpenStudy (anonymous):

Can you factor it a bit?

OpenStudy (anonymous):

FyYz(Xz+1)+FxXz(Zy+1)+FzZy(Yz+1)+ZyYzXz+1

OpenStudy (anonymous):

I wonder if \[ F_yy_z=-F_xx_z-F_z \]Wouldn't help cancel stuff out.

OpenStudy (anonymous):

Wait!

OpenStudy (anonymous):

You could expand out \(z_yy_zx_z\)

OpenStudy (anonymous):

Also, why isn't there a negative sign? Where did it go?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio So you'd say \[ z_x=-\frac{F_x+F_yy_x}{F_z} \] \(\color{blue}{\text{End of Quote}}\) Let's start from here.

OpenStudy (anonymous):

it is negative

OpenStudy (anonymous):

The whole thing is i believe

OpenStudy (anonymous):

I think you're on to something swapping things out with the original equations